# Boosting HDR Image Reconstruction via Semantic Knowledge Transfer

### v1

@@ 3.2 Semantic Priors Guided Reconstruction Model
3.3 Semantic Knowledge Transfer Scheme
@@
Then, the MSE loss between the two histograms is used as the histogram loss:
$\mathcal{L}_{\text{hist}}=\frac{1}{C}\sum_{c=1}^{C}\sum_{i=0}^{N-1}\left(H_{c,i}^{\text{S}}-H_{c,i}^{\text{T}}\right)^{2},$ (9)
where $H^{\text{S}}$ and $H^{\text{T}}$ represent the color histograms of the student and teacher model’s outputs, respectively. Building on Eq. (9), the proposed $\mathcal{L}_{\text{color}}$ is formulated as:
$\mathcal{L}_{\text{color}}=\frac{1}{N}\sum_{i=0}^{N-1}\text{Hist}\left(p^{i}_{m}\odot\mathcal{T}(H_{\text{in}}),p^{i}_{m}\odot\widehat{H_{\text{s}}}\right),$ (10)
where $\odot$ denotes element-wise multiplication, $p^{i}_{m}$ is the $i$-th channel of the segmentation maps derived from SKB, and $\text{Hist}(\cdot)$ is the histogram loss computed over each region.
@@
3.4 Semantic Knowledge Alignment Module
@@
The motivation behind SKAM is twofold. First, since different domains/formats share the same semantic information, knowledge transfer of feature-level can be achieved within a unified latent space. Second, inspired by the masked reconstruction [15], we aim to perform feature-level distillation in a more gradual manner, rather than directly minimizing feature distances. Specifically, as shown in Fig. 3, SKAM comprises two encoders, a masked feature mixer, and a decoder. At the $k$-th alignment stage, feature maps from the ORM ($F^{k}_{S}$) and SPGRM ($F^{k}_{T}$) are processed by their respective encoders, mapping them into a unified latent space to produce latent representations $\hat{F^{k}_{S}}$ and $\hat{F^{k}_{T}}$. These representations are then blended using a pair of randomly generated complementary masks. The decoder subsequently transforms the combined features into an enhanced feature $F^{k}_{E}$ in the same space as $F^{k}_{T}$, expressed as:
@@
3.5 Training and Inference
In the training stage, the SPGRM is exclusively supervised using a domain-transferred ground truth and optimized through an $\mathcal{L}_{1}$ loss, denoted as $L_{\text{spg}}$:
@@ 3.5 Training and Inference
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters empirically set to 1e-3 and 100. During the inference stage, only ORM participates in the inference, without requiring SPGRM.
Methods Tel’s dataset [28] Kalantari’s dataset [12]
PSNR-$\mu$uparrow PSNR-L $\uparrow$ SSIM-$\mu$uparrow SSIM-L $\uparrow$ HV2 $\uparrow$ PSNR-$\mu$uparrow PSNR-L $\uparrow$ SSIM-$\mu$uparrow SSIM-L $\uparrow$ HV2 $\uparrow$
NHDRR [38] 36.68 39.61 0.9590 0.9853 65.41 42.41 41.43 0.9877 0.9857 61.21
DHDR [12] 40.05 43.37 0.9794 0.9924 67.09 42.67 41.23 0.9888 0.9846 65.05
HDRGAN [24] 41.87 45.23 0.9832 0.9939 67.03 43.92 41.57 0.9905 0.9865 65.45
DiffHDR [40] 42.19 45.66 0.9843 0.9945 69.88 44.11 41.73 0.9911 0.9885 65.52
HyHDR [39] - - - - - 44.64 42.47 0.9915 0.9894 66.05
SCTNet [28] 42.48 47.46 0.9849 0.9952 70.66 44.43 42.21 0.9918 0.9891 66.64
LF-Diff [11] - - - - - 44.76 42.59 0.9919 0.9906 66.54
AHDR [37] 41.92 45.20 0.9836 0.9942 67.09 43.62 41.03 0.9900 0.9862 64.60
AHDR+ 43.18 46.97 0.9874 0.9955 70.39 44.24 41.54 0.9910 0.9880 65.84
+1.26 +1.77 +0.38 +0.13 +3.30 +0.62 +0.51 +0.10 +0.18 +1.24
HDR-Trans [22] 42.32 46.38 0.9843 0.9948 69.23 44.24 42.17 0.9916 0.9890 66.03
HDR-Trans+ 43.38 47.85 0.9879 0.9959 70.94 44.68 42.39 0.9918 0.9897 66.67
+1.06 +1.47 +0.36 +0.11 +1.71 +0.44 +0.22 +0.02 +0.07 +0.64
TMRNet [49] 42.72 45.98 0.9867 0.9953 69.13 43.92 42.09 0.9907 0.9891 65.32
TMRNet+ 43.15 47.19 0.9874 0.9961 70.59 44.49 42.39 0.9915 0.9895 66.04
+0.43 +1.21 +0.07 +0.09 +1.46 +0.57 +0.30 +0.08 +0.04 +0.72
Table 1: The evaluation results on Tel’s and Kalantari’s datasets. The improvements brought by our framework to the baseline models are highlighted in red, with the SSIM improvement displayed as 100$\times$ the actual value.
4 Experiments
4.1 Experimental Settings
@@
Datasets. We evaluate the proposed framework on several datasets from various scenes, including Kalantari’s dataset [12], Tel’s dataset [28], and Zhang’s dataset [49]. Kalantari’s dataset is a real captured dataset comprising 74 samples for training and 15 samples for testing, focusing primarily on daytime dynamic scenes. Zhang’s dataset primarily focuses on low-light environments and includes two subsets: BracketlRE and BracketlRE+. BracketlRE contains LDR image sequences with various degradations, while BracketlRE+ extends this by including additional 4$\times$ super-resolution tasks. Each subset contains 1,045 pairs used for training, and the remaining 290 pairs used for testing. In contrast, Tel’s dataset offers a more balanced distribution of scenes and motions, with samples collected under diverse lighting conditions and encompassing a variety of motion types in equal proportions, with 108 samples for training and 36 for testing. It is worth noting that Kalantari’s and Tel’s datasets are in sRGB format, while Zhang’s dataset uses RAW format.
Evaluation Metrics. To ensure a fair quantitative comparison, we adhere to the original metrics settings for each dataset. For Zhang’s dataset in RAW format, we use PSNR-$\mu$, SSIM-$\mu$, and LPIPS [46] as evaluation metrics. For Kalantari’s and Tel’s datasets in sRGB format, we use PSNR-L, PSNR-$\mu$, SSIM-L, SSIM-$\mu$, and HDR-VDP-2 [23]. The subscripts $\mu$ and L indicate that these metrics are computed in the tonemapped and linear domains.
Compared Methods. To verify the effectiveness of our designs, we conduct experiments on four datasets using several well-established HDR reconstruction models. These include AHDR [37], which is representative of CNN-based methods, and HDR-Trans [22], which exemplifies Transformer-based methods. Additionally, we include TMRNet [49], which is currently the SOTA method for HDR image reconstruction and enhancement. We further evaluate our method against a rich collection of SOTA techniques [5, 13, 28, 25, 38, 12, 4]. These models and datasets are carefully selected to cover a range of HDR reconstruction methodologies and scenarios, thereby demonstrating the versatility of our framework.
@@
4.2 Quantitative Evaluation
Quantitative results on RAW format dataset. The evaluation results on Zhang’s dataset are shown in Tab. 2. By introducing the semantic prior of self-distillation learning at the content and color levels, our method provides an average improvement of 2.14/1.98 dB in PSNR on the BracketIRE/BracketIRE+ datasets, respectively. Furthermore, TMRNet+ achieves PSNR values of 39.71/30.61 dB on the BracketIRE/BracketIRE+ datasets, setting a new state-of-the-art benchmark. Furthermore, our framework significantly enhances the SSIM index, underscoring its superior ability to restore the image dynamic range while maintaining structural integrity and fine details. Additionally, the notable increase in LPIPS scores, facilitated by our framework, demonstrates a closer alignment with human perceptual judgment. This is achieved through the innovative incorporation of semantic priors into our design.
Quantitative results on sRGB format dataset. The evaluation results on Tel’s and Kalantari’s datasets are shown in Tab. 1. We can observe that our framework effectively boosts the performance of baseline methods, providing an average improvement of 0.91/0.54 dB in PSNR on the Kalantari/Tel’s datasets, respectively. The PSNR improvement on the Kalantari’s dataset is only half that of the Tel’s dataset. This is because, compared to the Tel’s dataset, the Kalantari’s dataset primarily focuses on daytime dynamic scenes, where input images typically do not contain degradations, which are the main factors addressed by semantic priors. However, our proposed framework consistently improves the baseline’s performance. Through consistent and stable performance enhancements across multiple scene datasets, our framework demonstrates its effectiveness and flexibility.
@@ 4.2 Quantitative Evaluation
4.3 Qualitative Evaluation
Qualitative results on bright scenes. Reconstructing HDR images in dynamic bright scenes often causes ghosting artifacts when essential content (e.g., content, details) for overexposed areas is missing due to object or camera movement. Fig. 4 showcases qualitative results for bright scenes from the two datasets. Our framework is applied to multiple baseline models and compared with various state-of-the-art approaches using test data from the Kalantari’s (top row) and Tel’s (bottom row) datasets, which include challenging samples characterized by saturated backgrounds and foreground motions. As shown in the top row of Fig. 4, all baseline results exhibit noticeable artifacts in the step regions, whereas our proposed framework significantly suppresses ghosting phenomena. In the bottom row, all improved methods except AHDR+ eliminate ghosting, but AHDR+ still shows a significant reduction in ghosting compared to the baselines. This demonstrates the effectiveness of our framework in removing or suppressing ghosting caused by missing information.
Qualitative results on low-light scenes. Captured LDR images often suffer from noticeable degradation in low-light scenes, which significantly affects the final reconstruction quality. Fig. 5 showcases qualitative results from the BracketIRE (left) and BracketIRE+ (right) datasets. As shown in the left part, the LDR images contain noticeable degradations, and all methods exhibit unclear details. Our proposed framework significantly enhances the model’s detail representation, reducing the impact of image degradations. For example, the tigers and stars within the blue boxes exhibit clearer textures. Furthermore, compared to BracketIRE, the LDR images in BracketIRE+ have lower resolution, posing a challenge for reconstructing high-quality HDR images. As shown in the right part, all baseline models almost fail to restore realistic textures and show significant blurring. In contrast, our proposed framework significantly improves the imaging quality of the baseline models, recovering the scene’s outlines and details more completely.

### v2

@@ III-B Semantic Priors Guided Reconstruction Model
III-C Semantic Knowledge Transfer Scheme
@@
To obtain a proper probability distribution, we normalize the histogram by dividing each $H_{c,i}$ by the sum of $H_{c}$ for numerical stability. Then, the mean squared error (MSE) between the $\mathrm{CDF}_{c,k}=\sum_{i=0}^{k}H_{c,i},k\in\{0,\ldots,N-1\}$ of the two histograms, is used as the histogram loss:
$\mathcal{L}_{\text{hist }}=\frac{1}{C}\sum_{c=1}^{C}\sum_{k=0}^{N-1}\left(\mathrm{CDF}_{c,k}^{\mathrm{S}}-\mathrm{CDF}_{c,k}^{\mathrm{T}}\right)^{2},$ (9)
where $\mathrm{CDF}_{c,k}^{\mathrm{S}}$ and $\mathrm{CDF}_{c,k}^{\mathrm{T}}$ represent the CDFs derived from the color histograms of the output of the student and teacher model, respectively. Building on Eq. (9), the proposed $\mathcal{L}_{\text{color}}$ is formulated as:
[…]
where $\odot$ denotes element-wise multiplication, $p^{i}_{m}$ is the $i$-th channel of the binary segmentation maps derived from SKB, and $\text{Hist}(\cdot)$ is the histogram loss computed over each region.
@@
III-D Semantic Knowledge Alignment Module
@@
The motivation behind SKAM is twofold. First, since different domains/formats share the same semantic information, knowledge transfer of feature-level can be achieved within a unified latent space. Second, inspired by the masked reconstruction [23], we aim to perform feature-level distillation in a more gradual manner, rather than directly minimizing feature distances. Specifically, as shown in Fig. 3, SKAM is composed of two encoders, one masked feature mixer, and one decoder. The encoders and the decoder possess the same architectural design, each consisting of two convolutional layers. At the $k$-th alignment stage, feature maps from the ORM ($F^{k}_{S}$) and SPGRM ($F^{k}_{T}$) are processed by their respective encoders, mapping them into a unified latent space to produce latent representations $\hat{F^{k}_{S}}$ and $\hat{F^{k}_{T}}$. These representations are then fused using a set of two randomly generated binary complementary masks, with a 0.5 probability for each value in the masks. The decoder subsequently transforms the combined features into an enhanced feature $F^{k}_{E}$ in the same space as $F^{k}_{T}$, expressed as:
@@
Figure 4: The process of the training and inference stages.
III-E Training and Inference
As shown in Fig. 4, in the training stage, the SPGRM is exclusively supervised using a domain-transferred ground truth and optimized through an $\mathcal{L}_{1}$ loss, denoted as $L_{\text{spg}}$:
@@ III-E Training and Inference
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters empirically set to 1e-3 and 100. To tackle the variability in the number of segmentation instances output by SAM-based models, we fix the number of output instance masks to a constant value $K=50$ for every training sample. Notably, only non-zero (valid) instance masks are involved in the calculation of $\mathcal{L}_{color}$. During the inference stage, only ORM participates in the inference, without requiring SPGRM.
[…]
Methods Tel et al.’s dataset [38]
PSNR-$\mu$ $\uparrow$ SSIM-$\mu$ $\uparrow$ PSNR-L $\uparrow$ SSIM-L $\uparrow$ PSNR-PU $\uparrow$ SSIM-PU $\uparrow$ HDR-VDP2 $\uparrow$
NHDRR [53] 36.68 0.9590 39.61 0.9853 37.06 0.9777 67.52
DHDR [17] 40.05 0.9794 43.37 0.9924 40.46 0.9889 68.08
HDRGAN [35] 41.87 0.9832 45.23 0.9939 42.13 0.9908 70.14
DiffHDR [50] 42.19 0.9843 45.66 0.9945 42.51 0.9919 71.52
SMHDR [33] 43.41 0.9883 47.46 0.9952 43.79 0.9941 72.76
AFUNet [25] 43.31 0.9876 47.83 0.9959 43.72 0.9945 73.22
SCTNet [38] 42.48 0.9849 47.46 0.9952 42.79 0.9924 71.99
AHDR [49] 41.92 0.9836 45.20 0.9942 42.29 0.9919 70.19
AHDR+ 43.18 0.9874 46.97 0.9955 43.61 0.9937 72.34
+1.26 +0.38 +1.77 +0.13 +1.32 +0.18 +2.15
HDR-Trans [31] 42.32 0.9843 46.38 0.9948 42.64 0.9920 70.99
HDR-Trans+ 43.38 0.9879 47.85 0.9959 43.84 0.9943 72.88
+1.06 +0.36 +1.47 +0.11 +1.20 +0.23 +1.89
TMRNet [62] 42.72 0.9867 45.98 0.9953 43.23 0.9936 72.15
TMRNet+ 43.15 0.9874 47.19 0.9961 43.47 0.9939 72.22
+0.43 +0.07 +1.21 +0.08 +0.24 +0.03 +0.07
[…]
Methods Kalantari et al.’s dataset [17]
PSNR-$\mu$ $\uparrow$ SSIM-$\mu$ $\uparrow$ PSNR-L $\uparrow$ SSIM-L $\uparrow$ PSNR-PU $\uparrow$ SSIM-PU $\uparrow$ HDR-VDP2 $\uparrow$
NHDRR [53] 42.41 0.9877 41.43 0.9857 42.97 0.9855 66.69
DHDR [17] 42.67 0.9888 41.23 0.9846 41.83 0.9832 70.58
HDRGAN [35] 43.92 0.9905 41.57 0.9865 44.03 0.9851 71.13
DiffHDR [50] 44.11 0.9911 41.73 0.9885 43.81 0.9914 71.19
HyHDR [48] 44.64 0.9915 42.47 0.9894 45.25 0.9928 71.96
SMHDR [33] 44.69 0.9920 42.61 0.9894 45.30 0.9932 71.94
AFUNet [25] 44.91 0.9923 42.59 0.9906 45.57 0.9935 71.98
SCTNet [38] 44.43 0.9918 42.21 0.9891 44.27 0.9927 71.81
LF-Diff [16] 44.76 0.9919 42.59 0.9906 45.28 0.9930 72.11
AHDR [49] 43.62 0.9900 41.03 0.9862 42.93 0.9849 71.08
AHDR+ 44.24 0.9910 41.54 0.9880 44.99 0.9925 71.24
+0.62 +0.10 +0.51 +0.18 +2.06 +0.76 +0.16
HDR-Trans [31] 44.24 0.9916 42.17 0.9890 44.23 0.9924 70.22
HDR-Trans+ 44.68 0.9918 42.39 0.9897 45.27 0.9928 71.94
+0.44 +0.02 +0.22 +0.07 +1.04 +0.04 +1.72
TMRNet [62] 43.92 0.9907 42.09 0.9891 44.41 0.9918 71.45
TMRNet+ 44.49 0.9915 42.39 0.9895 45.11 0.9928 71.79
+0.57 +0.08 +0.30 +0.04 +0.70 +0.10 +0.34
IV Experiments
IV-A Experimental Settings
@@
Datasets. We evaluate the proposed framework on several datasets spanning diverse scenes, including Kalantari’s dataset [17], Tel et al.’s dataset [38], and Zhang et al.’s dataset [62]. These datasets cover a wide range of scene conditions, encompassing variations in lighting, object motion, and camera motion. Kalantari et al.’s dataset is a real-captured dataset consisting of 74 training samples and 15 testing samples, focusing primarily on daytime dynamic scenes. Zhang et al.’s dataset, by contrast, focuses on low-light environments and includes two subsets: BracketIRE and BracketIRE+. BracketIRE contains SDR image sequences with various degradations, while BracketIRE+ extends this by incorporating an additional 4× super-resolution task. Each subset comprises 1,045 pairs for training and 290 pairs for testing. Tel et al.’s dataset offers a more balanced distribution of scenes and motions, with samples collected under diverse lighting conditions and encompassing various motion types in equal proportions; it includes 108 training samples and 36 testing samples. To further assess the model’s ability to generalize, we also make use of Sen et al.’s dataset [36] and Tursun et al.’s dataset [40]; these datasets are exclusively used for qualitative evaluation, given that they do not provide corresponding ground truths. Notably, Kalantari’s and Tel’s datasets are in sRGB format, with three SDR images per scene as input. In comparison, Zhang’s dataset uses RAW format, with five SDR images per scene as input.
@@
Evaluation Metrics. For Zhang et al.’s dataset in RAW format, we follow the evaluation protocol established in the original BracketIRE dataset [62]. Specifically, RAW images are first converted to linear RGB using a defined ISP pipeline, then tone-mapped to 16-bit SDR for computing PSNR, SSIM, and LPIPS [60]. All metrics are evaluated against the corresponding SDR ground truth. For Kalantari et al.’s and Tel et al.’s datasets in sRGB format, we adopt the same metrics as prior work [49, 31, 38]: PSNR-L, PSNR-$\mu$, SSIM-L, SSIM-$\mu$, and HDR-VDP2 [32] to enable fair and intuitive comparison. The subscripts L and $\mu$ denote evaluation in the linear and $\mu$-law tone-mapped domains, respectively. We calculate HDR-VDP2 using linear input images with the “rgb-native” color encoding configuration. The data is treated as display-referred and scaled to a peak luminance of $L_{peak}=1000$ nits, with a viewing angle of $30^{\circ}$ and a viewing distance of 0.55 meters. In addition, we extend our evaluation suite with PSNR-PU and SSIM-PU, computed using the official PU21 implementation [3], which are perceptually optimized metrics specifically designed for HDR quality assessment.
Compared Methods. To verify the effectiveness of our designs, we conduct experiments on four datasets using several well-established HDR reconstruction models. These include AHDR [49], which is representative of CNN-based methods, and HDR-Trans [31], which exemplifies Transformer-based methods. Additionally, we include TMRNet [62], which is currently the SOTA method for HDR image reconstruction and enhancement. We further evaluate our method against a rich collection of SOTA techniques [53, 50, 33, 16, 25, 48, 38, 34, 53, 17]. Specifically, for Zhang et al.’s dataset [62], we additionally include two models: Burstformer [9] and BIPNet [8], as representative works in burst image restoration and enhancement. We also incorporate Kim et al.’s method [19], which is specifically designed for RAW-domain HDR tasks (e.g., demosaicing and deghosting) and thus aligns with the dataset’s RAW format. These models and datasets are carefully selected to cover a range of HDR reconstruction methodologies and scenarios, thereby demonstrating the versatility of our framework.
TABLE III: The evaluation results on Zhang’s dataset.
@@
IV-B Quantitative Evaluation
Quantitative results on RAW format dataset. The evaluation results on Zhang et al.’s dataset are shown in Tab. III. The benchmark matches the results of the NTIRE [63] version and includes official weights. By introducing the semantic prior of self-distillation learning at the content and color levels, our method provides an average improvement of 2.14/1.98 dB in PSNR on the BracketIRE/BracketIRE+ datasets, respectively. Furthermore, TMRNet+ achieves PSNR values of 39.71/30.61 dB on the BracketIRE/BracketIRE+ datasets, setting a new state-of-the-art benchmark. Furthermore, our framework significantly enhances the SSIM index, underscoring its superior ability to restore the image dynamic range while maintaining structural integrity and fine details. Additionally, the notable increase in LPIPS scores, facilitated by our framework, demonstrates a closer alignment with human perceptual judgment. This is achieved through the innovative incorporation of semantic priors into our design.
Quantitative results on sRGB format dataset. The evaluation results on Tel et al.’s and Kalantari et al.’s datasets are presented in Tab. I and Tab. II, respectively. It can be observed that our framework effectively enhances the performance of baseline methods, achieving an average improvement of 1.26 dB and 0.92 dB in PSNR-PU on Kalantari et al.’s and Tel et al.’s datasets, respectively. Notably, when applied to our proposed framework, TMRNet yields a relatively smaller performance gain compared to AHDR and HDR-Transformer. This phenomenon can be attributed to the fact that TMRNet is inherently designed for HDR reconstruction scenarios with degradations, whereas the latter two baseline models focus more on dynamic motion scenes. Consequently, the performance improvement brought by semantic prior injection is relatively limited for TMRNet. Nevertheless, our proposed framework consistently enhances the performance of all baseline models. By delivering consistent and stable performance improvements across multiple scene-specific datasets, our framework fully demonstrates its effectiveness and flexibility.
@@ IV-B Quantitative Evaluation
IV-C Qualitative Evaluation
Qualitative results on bright scenes. Reconstructing HDR images in dynamic bright scenes often causes ghosting artifacts when essential content (e.g., content, details) for overexposed areas is missing due to object or camera movement. Fig. 5 showcases qualitative results for bright scenes from the two datasets. Our framework is applied to multiple baseline models and compared with various state-of-the-art approaches using test data from the Kalantari et al.’s (top row) and Tel et al.’s (bottom row) datasets, which include challenging samples characterized by saturated backgrounds and foreground motions. As shown in the top row of Fig. 5, all baseline results exhibit noticeable artifacts in the step regions, whereas our proposed framework significantly suppresses ghosting phenomena. In the bottom row, all improved methods except AHDR+ eliminate ghosting, but AHDR+ still shows a significant reduction in ghosting compared to the baselines. This demonstrates the effectiveness of our framework in removing or suppressing ghosting caused by missing information.
Qualitative results on low-light scenes. Captured SDR images often suffer from noticeable degradation in low-light scenes, which significantly affects the final reconstruction quality. Fig. 6 showcases qualitative results from the BracketIRE (left) and BracketIRE+ (right) datasets. As shown in the left part, the SDR images contain noticeable degradations, and all methods exhibit unclear details. Our proposed framework significantly enhances the model’s detail representation, reducing the impact of image degradations. For example, the tigers and stars within the blue boxes exhibit clearer textures. Furthermore, compared to BracketIRE, the SDR images in BracketIRE+ have lower resolution, posing a challenge for reconstructing high-quality HDR images. As shown in the right part, all baseline models almost fail to restore realistic textures and show significant blurring. In contrast, our proposed framework significantly improves the imaging quality of the baseline models, recovering the scene’s outlines and details more completely.
Qualitative results on dataset without ground truth. To further validate the generalization capability of the proposed method, we additionally assess its performance using the datasets from Sen et al.[36] and Tursun et al.[40], neither of which provides ground truth, as shown in Fig. 7. On the Sen et al.’ dataset (left), severe overexposure and fine-scale textures cause AHDR [49], HDR-Trans [31], and TMRNet [62] to produce blown-out regions with texture loss or over-exposure. In contrast, when our framework is applied to distill semantic priors from FastSAM [64] and enforce color-consistency regularization, these models recover clearer details, stabilize local contrast, and yield more natural highlight roll-off. On the Tursun et al.’s dataset (right), extreme luminance gradients cause AHDR and TMRNet to generate ghosting near the backpack boundary, while HDR-Trans loses fine structure. After integrating our framework, the same baselines preserve sharp boundaries, maintain structural coherence, suppress ghosting, and reconstruct smooth and physically plausible luminance gradients. These results demonstrate that our framework generalizes well to challenging real-world scenes with complex exposure conditions and motion, consistently improving both visual quality and structural fidelity.
TABLE IV: Ablation studies of proposed framework.